# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### 13 Dec 2023 v1

@@ II-B File Structure
In the root directory MTO/ of MToP, there is a running script file mto.m and four folders, Algorithms/, Problems/, Metrics/, and GUI/. The structure of these folders is shown in Fig. 4. The algorithms folder Algorithms/ holds various categories folder, utility functions, and the algorithm base class Algorithm.m. The specific algorithm files, e.g., MFEA.m, MO_MFEA.m, GA.m, are stored under their corresponding classification folder. The utility functions folder Algorithms/Utils/ provides various public utility functions of generation operators, environmental selection approaches, constraint handling techniques, etc. The problems folder Problems/ contains various categories folder, base function folder, and the problem base class Problem.m. The specific problem files, e.g., WCCI20_MTSO1.m, CMT3.m, CEC19_MaTMO6.m, are stored under their corresponding classification folder. In particular, files for real-world application problems are placed in Real-World Application/ folder. The Problems/Base/ folder includes the public base functions for problems. The third folder Metrics/ in MToP contains all the result evaluation metrics and is with two subfolders, Metrics/Single-objective/ and Metrics/Multi-objective for single-objective and multi-objective optimization respectively. Specific metric files are Obj.m, Obj_MTS.m, IGD.m, etc. The last folder is the GUI/ which contains all the files used by the GUI of MToP. Among them, MTO_GUI.m is the main file of the GUI, and MTO_CMD.m provides the function of executing experiments by the command line.
@@ II-C1 Algorithm
TABLE VII: Properties and methods of algorithm base class.
@@
2    caption={Algorithm implementation example.},
@@
6classdef Algo_Example < Algorithm
@@
14  function run(Algo, Prob)
15    % initialize populations for all tasks
16    pops = Initialization(Algo, Prob, Individual);
@@
18    while Algo.notTerminated(Prob, pops) % main loop
@@
21        off = Generation(pops{t});
@@
All algorithms in MToP inherit from the algorithm base class Algorithm.m. The properties and methods of the algorithm base class are shown in Table VII. The base class provides all the functions that interact with the GUI, so the implemented algorithms only need to consider the evolutionary workflow itself. A simple example of algorithm implementation is shown in Listing LABEL:list:algo. The second line here is labels for the algorithms, inspired by PlatEMO [28], for the algorithm classification displayed in the GUI. The function run() of the class is for the execution of the algorithm, with its first argument Algo denoting the object itself and the second argument Prob denoting the problem object to be solved. At the beginning of the run() function, the population initialization is performed with the Initialization() function. MToP provides two population initialization functions, the multi-population method and the multifactorial method, which produce either multiple populations or single populations with skill factors respectively. Here the pops produced in line 11 are of cell data type containing multiple populations. Each individual in the pops is Individual object and contains properties including objective value Obj, constraint violation CV, and decision variables Dec. Then, the main loop is entered with the function notTerminated() implemented in the algorithm base class. The function notTerminated() provides data updates and generation counting methods. In the main loop, the functions of offspring generation Generation(), offspring evaluation Evaluation(), and environmental selection Selection() are performed separately for each task. For algorithms with the multifactorial framework, these functions are executed simultaneously for all tasks. The function Generation() acts on the decision variables Dec of offspring individuals and is implemented by specific algorithms. The evaluation function Evaluation() is implemented in the algorithm base class, which automatically calls the problem object Prob for offspring fitness evaluation, i.e., calculating the objective value Obj and constraint violation CV for all individuals. It also updates the best solution automatically for single-objective optimization. At the end of each loop, through the environmental selection function Selection(), the new populations pops are updated. MToP provides some universal Selection() functions, and this function can also be reimplemented by specific algorithms. After that, the code goes to the next loop and calls notTerminated() to record changes. The algorithm code pattern in MToP is adapted to all EAs for solving MTO problems.
@@ II-C2 Problem
TABLE VIII: Properties and methods of problem base class.
@@ 11    Prob = Prob@Problem(varargin);
14  end
15
16  function setTasks(Prob)
@@
30  end
31
32  function optimum = getOptimum(Prob)
@@
36  end
37end
All problems in MToP inherit from the problem base class Problem.m. The properties and methods of the Problem.m are shown in Table VIII. A simple example of problem implementation is presented in Listing LABEL:list:prob. The maximum number of function evaluations maxFE for the problem can be defined in the constructor of the class. The function setTasks() is for setting the properties and evaluation function for each optimization task. In this function, the number of tasks T of the problem is first defined followed by a detailed per-task setup. After that, decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub need to be set separately for each task. M and D for multiple tasks of the problem are denoted as vector type. Fnc, Lb and Ub are denoted as cell type. Note that the function Fnc for each task takes inputs of the decision variable Dec and returns objective values Obj and the constraint violation value CV. Additionally, for multi-objective optimization, the external function getOptimum(), which obtains optimal solutions (i.e., reference points), needs to be defined. With these problem code patterns, all problems that are completely defined can be solved using the corresponding type of algorithm in MToP.
II-C3 Metric
@@
In contrast to algorithm and problem class implementation, metric codes in MToP are recognized as the function type. A simple example of metric implementation is shown in Listing LABEL:list:metric. The input parameter of the function is MTOData generated by the experimental execution. The function returns result, which contains properties of Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData. The property Metric can be either ‘Min’ or ‘Max’, indicating that a smaller or larger metric value is better. RowName, ColumnName, and TableData are used to return the metric results for the GUI table. ConvergeData and ParetoData are used to return metric convergence results for convergence plots and non-dominated solutions for Pareto plots respectively. Note that specific metric calculation functions are not covered here for the simplicity of this example.
@@ III Performing Experiments with MToP
Running MToP requires first executing the script file mto.m in the root directory. After that, the MToP GUI is shown as presented in Fig. 5.
@@ III-A Testing Algorithms and Problems
As shown in Fig. 5 (a), the test module is divided into left, center, and right parts. The left part is for the selection and setup of algorithms and problems. After selecting the corresponding labels of Task, Objective, and Special drop-down buttons, the Algorithm and Problem drop-down buttons update their contents. Selecting specific algorithms and problems displays the parameters they contain in the edit box, which can be also clicked on and set. After parameters are updated, new parameter values are fed back into the algorithm and problem objects through the setParameter() method. The parameters of the algorithms generally contain evolutionary operator parameters and their customized parameters. The problem parameters generally contain population size N, maximum function evaluations maxFE, and their customized parameters. The center part of the MToP GUI can execute the test experiment and display the graphic results. After executing the procedure of the algorithm on the problem, the metric results of the problem are displayed in the right part of the test module.
@@ III-B Executing Comparative Experiments
The GUI of the experiment module is shown in Fig. 5 (b). It is divided into a two-column experimental setup part on the left and a results display part on the right. The first column contains the basic settings to perform experiments. The text box No. of Runs indicates the number of independent repeat runs for each algorithm on each problem. The text box No. of Results denotes the number of results saved by the evolutionary process of each algorithm, which is uniformly sampled according to the number of generations. The Parallel button controls whether experiments are executed in parallel processing. Parallel execution invokes multiple threads to run simultaneously multiple independent runs. The drop-down buttons Task, Objective, and Special serve the same purpose as the test module, filtering algorithms and problems based on their pre-defined labels. After selecting the algorithm and problem in the two boxes on the bottom side of the first column, it can be added to the preparation area of the second column by clicking the Add button. The names and parameters of the added algorithms and problems can be edited by clicking on them. After these are ready, click on the green Start button and MToP starts executing the experiment. It is possible to click on the buttons Pause and Stop to pause or stop the experiment during the execution.
When the experiment has been completed, the content of the table is available in the results area on the right part of the experiment module. The original data of the experiment can be read and saved by Load Data and Save Data buttons. In the right part, The buttons in the first function row operate on the experimental data. The first button Save Table in the line can export the results of the table to xlsx, csv, and tex files. The second and third Converge buttons control convergence y-lim types and draw convergence plots based on the results selected in the table. The fourth button Pareto draws Pareto front and non-dominated solutions based on the selected results. The fifth edit box controls the type of result formatting of the table. The drop-down box right of the edit box is for selecting performance metrics. The drop-down boxes following the metrics are the result display type, statistical test method, base algorithm, and result highlighting type, respectively. After the statistical test method is selected, the calculated statistical results, denoted by “+/-/=” followed by each metric result, can be also displayed in the table. MToP provides two non-parametric methods, the Wilcoxon signed rank test and the Wilcoxon rank-sum test, for statistical tests. The base algorithm drop-down button controls the base algorithm in the statistical test. The highlight button highlights the best and worst algorithm’s metrics results for each problem or task.
@@ III-C Executing via Alternative Command Line
In addition to GUI execution, MToP provides alternative options for running experiments via the command line. After inputting the algorithm list, problem list and other parameters into the script file mto.m, the experiments can be run directly in the command line and the data MTOData.mat are returned after running completed. Similarly, inputting the data into the metrics function yields the calculated metric results.
@@ IV Extending MToP
New algorithms, problems and performance metrics can be easily added in MToP. Users can create and implement code files following the code pattern described in Section II-C.
@@ IV-A Adding New Algorithms
The first step in adding an algorithm to MToP is to create a MATLAB class file that inherits the base class Algorithm.m in the Algorithms/ folder. The algorithmic framework can then be implemented following the code pattern as in Listing LABEL:list:algo. Users can use the population initialization function Initialization(), offspring generation function Generation(), and environmental selection function Selection() available in MToP, or of course reimplement these functions according to their requirements. In addition, users who want to create customized parameters that can be clicked and modified in the MToP GUI require an additional implementation of getParameter() and setParameter() API in the new algorithm class file.

### 09 Apr 2024 v2

@@ II-B File Structure
In the root directory MTO/ of MToP, you’ll find a script file named mto.m and four folders: Algorithms/, Problems/, Metrics/, and GUI/. The structure of these folders is depicted in Fig. 6.
The Algorithms/ folder contains various categories of algorithms, utility functions, and the algorithm base class Algorithm.m. Specific algorithm files such as MFEA.m, MO_MFEA.m, and GA.m are stored within their respective classification folders. The Algorithms/Utils/ folder provides public utility functions for generation operators, environmental selection approaches, constraint handling techniques, etc.
The Problems/ folder houses various categories of problems, the base function folder, and the problem base class Problem.m. Specific problem files like WCCI20_MTSO1.m, CMT3.m, and CEC19_MaTMO6.m are stored under their corresponding classification folders. Real-world application problem files are located in the Real-World Application/ folder, while the Problems/Base/ folder contains public base functions for problems.
The Metrics/ folder contains all result evaluation metrics, organized into subfolders for single-objective and multi-objective optimization (Metrics/Single-objective/ and Metrics/Multi-objective). Specific metric files include Obj.m, Obj_MTS.m, and IGD.m, among others.
Lastly, the GUI/ folder houses all files used by the GUI of MToP. Among them, MTO_GUI.m serves as the main file of the GUI, while MTO_CMD.m provides functionality for executing experiments via the command line.
@@ II-C1 Algorithm
TABLE III: Properties and methods of algorithm base class.
@@
2    caption={Algorithm implementation example for multi-population EA.},
@@
6classdef Algo_Example1_MP < Algorithm
@@
14function run(Algo, Prob)
15% initialize multiple populations for all tasks
16pops = Initialization(Algo, Prob, Individual);
@@
18while Algo.notTerminated(Prob, pops) % main loop
@@
21    off(index1) = Generation(pops{t});
22    % knowledge transfer
23    off(index2) = KnowledgeTransfer(pops)
@@
29end
30end
1    label={list:algo2},
2    caption={Algorithm implementation example for multifactorial EA.},
3    captionpos=b,
4    float=htbp
5]
6classdef Algo_Example2_MF < Algorithm
7% <Multi-task> <Multi-objective> <None/Constrained>
8
9properties % set parameter
10  para_example = 0.2;
11end
12
13methods
14function run(Algo, Prob)
15% initialize populations for multifactorial evolution
16population=Initialization_MF(Algo, Prob, Individual);
17
18while Algo.notTerminated(Prob, pops) % main loop
19  % generate new offspring with random mating
20  off = Generation(population);
21  % evaluate based on offspring skill factor
22  for t = 1:Prob.T
23    idx = [off.MFFactor] == t;
24    off(idx) = Algo.Evaluation(off(idx), Prob, t);
@@
All algorithms within MToP inherit from the algorithm base class Algorithm.m. The properties and methods of this base class are detailed in Table III. This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself. The MTEAs implemented in MToP are outlined in Table IV. Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems. Specific algorithms belonging to this category are listed in Table V.
Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively. The algorithm labels provided in the second line serve as identifiers for classification within the GUI. The run() function within each class orchestrates algorithm execution, with Algo representing the object itself and Prob denoting the problem object to be solved. Population initialization occurs at the outset of the run() function, utilizing either the multi-population or multifactorial method provided by MToP.
In Listing LABEL:list:algo, the variable pops is of cell data type, housing multiple populations as popspopulation1, population2, population3…. Each individual within pops is an Individual object, containing properties such as objective value (Obj), constraint violation (CV), and decision variables (Dec). Contrastingly, Listing LABEL:list:algo2 utilizes a list data type for the population variable, directly containing individuals as population(individual1:factor1, individual2:factor1, individual3:factor2…), along with an additional property (MFFactor).
Subsequently, the primary loop commences with the invocation of the notTerminated() function, a component of the algorithm base class. This function oversees data updates and generation counting within the loop. During the main loop, distinct operations are carried out for offspring generation Generation(), offspring evaluation Evaluation(), and environmental selection Selection().
In the context of knowledge transfer, the multi-population algorithm requires the implementation of the KnowledgeTransfer() function to acquire knowledge from other tasks. Conversely, the multifactorial approach achieves knowledge transfer through random mating within the Generation() function. The Generation() function operates on the decision variables (Dec) of offspring individuals and is tailored to specific algorithms. Meanwhile, the Evaluation() function, embedded within the algorithm base class, automatically engages the problem object (Prob) for offspring fitness evaluation. This entails calculating the objective value (Obj) and constraint violation (CV) for all individuals. Following each loop iteration, the environmental selection function Selection() is invoked to update the new population. While MToP offers universal Selection() functions, specific algorithms also have the option to reimplement this function. Subsequently, the code progresses to the next loop and invokes notTerminated() to document changes. The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.
TABLE IV: Multitask evolutionary algorithms included in MToP (current version).
ASCMFDE [53] 2021 TEVC Single Multi Multifactorial Yes Intertask transfer in aligned subspace for MFDE
MKTDE [56] 2022 TEVC Single Multi Multi-population Yes Meta-knowledge transfer-based multitask DE
MTEA-AD [23] 2022 TEVC Single Multi/Many Multi-population Yes MTEA with adaptive knowledge transfer via anomaly detection
MTES-KG [59] 2023 TEVC Single Multi/Many Multi-population Yes Multitask ES with knowledge-guided external sampling
TRADE [60] 2023 TCYB Single Many Multi-population Yes Transferable adaptive parameter DE for many-task optimization
MTEA-SaO [61] 2023 INS Single/Multi Multi/Many Multi-population Yes Adaptive solver multitask framework with implicit knowledge transfer
KR-MTEA [62] 2023 INS Single/Multi Multi/Many Multi-population Yes Adaptive multi-task EA based on knowledge reuse
MO-MFEA [42] 2017 TCYB Multi Multi Multifactorial Yes Multiobjective MFEA with non-dominated sort and crowding distance
EMT-PD [64] 2021 TSMC-S Multi Multi Multifactorial Yes Two-stage adaptive knowledge transfer based on population distribution
MTDE-MKTA [69] 2024 TEVC Multi Multi Multi-population Yes Multitask DE with multiple knowledge types and transfer adaptation
TABLE V: Single-task evolutionary algorithms for multitask optimization included in MToP (current version).
IPOP-CMA-ES [85] 2005 CEC Single Evolution Strategy Yes Restart CMA-ES with increasing population size
sep-CMA-ES [86] 2008 PPSN Single Evolution Strategy Yes Separable CMA-ES for large-scale optimization
C2oDE [95] 2019 TSMC-S Single Constraint Handling Yes Composite DE for constrained optimization
DeCODE [96] 2021 TSMC-S Single Constraint Handling Yes Decomposition-based multiobjective approach for constrained optimization
CEDE-DRL [98] 2023 SWEC Single Constraint Handling Yes Deep reinforcement learning assisted co-evolutionary DE
CCEF-ECHT [99] 2024 TSMC-S Single Constraint Handling Yes Competitive and cooperative ensemble of constraint handling techniques
@@ II-C2 Problem
TABLE VI: Properties and methods of problem base class.
@@ 11  Prob = Prob@Problem(varargin);
14end
15function setTasks(Prob)
@@
29end
30function optimum = getOptimum(Prob)
@@
34end
35end
All problems within MToP inherit from the problem base class Problem.m. The properties and methods of Problem.m are detailed in Table VI. As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach [19]. For task $k$ with upper and lower bounds $L_{k}$ and $U_{k}$, respectively, and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is linearly mapped as follows:
$\bm{x}^{\prime}=\frac{\bm{x}-L_{k}}{U_{k}-L_{k}}.$ (2)
The dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
A simple problem implementation example is illustrated in Listing LABEL:list:prob. The maximum number of function evaluations, denoted as maxFE, for the problem can be specified within the class constructor. The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task. Within this function, the number of tasks T is first defined, followed by a detailed setup for each task. Subsequently, parameters such as decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub are set individually for each task. For problems involving multiple tasks, M and D are represented as vector type, while Fnc, Lb, and Ub are represented as cell type. Notably, the function Fnc for each task takes the decision variable Dec as input and returns objective values Obj along with the constraint violation value CV. Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP. The existing problems available in MToP are enumerated in Table VII. In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
TABLE VII: Multitask benchmark and real-world optimization problems included in MToP (current version).
Problem Case Task Dimension Objective Special Description
SOPM [69] 2 3 30 Multi Constrained Synchronous optimal pulse-width modulation as constrained multiobjective multitask problem
MO-OPF [69] 4 2 24/33 Multi Constrained Multiobjective optimal power flow as constrained multiobjective multitask problem
MO-SCP [61] 1 Any 60-120 Multi Competitive Multiobjective sensor coverage problem as competitive multiobjective multitask problem
II-C3 Performance Metric
@@
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is MTOData, which is generated during experimental execution. The function returns result, comprising properties such as Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData. The Metric property can take values of either Min’ or Max’, indicating whether a smaller or larger metric value is preferable. RowName, ColumnName, and TableData are utilized to present metric results in the GUI table. On the other hand, ConvergeData and ParetoData are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table VIII.
TABLE VIII: Metrics included in MToP (current version).
Metric Objective Task Description
Obj Single All Objective value for each task
Obj (MTS) Single Multi Multitask score of Obj for all tasks
Obj (AV) Single Multi Average Obj for all tasks
Obj (CMT) Single Multi Competitive multitask Obj for all tasks
Obj (NBR) Single Multi Number of best Obj result for all tasks
CV Single All Constraint violation for each task
FR Single All Feasible rate for each task
HV Multi All Hypervolume for each task
HV (MTS) Multi Multi Multitask score of HV for all tasks
HV (CMT) Multi Multi Competitive multitask HV for all tasks
IGD Multi All Inverted generational distance for each task
IGD+ Multi All Improved IGD plus for each task
IGD+ (MTS) Multi Multi Multitask score of IGD+ for all tasks
IGD+ (CMT) Multi Multi Competitive multitask IGD+ for all tasks
Run Time All All Algorithm running time for all tasks
@@ III Performing Experiments with MToP
To launch MToP, start by running the script file mto.m located in the root directory. This action will initialize the MToP GUI interface, which is illustrated in Fig. 7.
@@ III-A Testing Algorithms and Problems
As depicted in Fig. 7 (a), the test module interface is structured into left, center, and right sections. The left section facilitates the selection and configuration of algorithms and problems. Upon choosing the desired labels such as Task, Objective, and Special from the drop-down buttons, the options available for Algorithm and Problem are updated accordingly. Selecting specific algorithms and problems reveals their parameters in the edit box, which can be edited by clicking on them. Once parameters are adjusted, the new values are applied to the algorithm and problem objects through the setParameter() method. Algorithm parameters typically include evolutionary operator settings and other customizable parameters, while problem parameters usually consist of population size N, maximum function evaluations maxFE, and other specific parameters.
In the center section of the MToP GUI, users can execute the test experiment and visualize the results graphically. Upon completion of the algorithm execution on the chosen problem, the metric results of the problem are displayed in the right section of the test module.
@@ III-B Executing Comparative Experiments
The GUI of the experiment module is displayed in Fig. 7 (b), featuring a three-column layout. The left and middle columns comprise the experimental setup, while the right column displays the results.
In the first column, users can configure the basic experiment settings. The No. of Runs textbox specifies the number of independent repeat runs for each algorithm on each problem. The No. of Results textbox determines the number of results saved during the evolutionary process of each algorithm, sampled uniformly according to the number of generations. The Parallel button toggles parallel processing, enabling multiple threads to run independent experiments simultaneously. The drop-down buttons Task, Objective, and Special filter algorithms and problems based on predefined labels, similar to the test module. Users can add selected algorithms and problems by clicking the Add button after choosing them from the respective drop-down boxes at the bottom.
In the second column, the names and parameters of the added algorithms and problems can be edited by clicking on them. Once all configurations are set, clicking the green Start button initiates the experiment execution. Users can also pause or stop the experiment using the Pause and Stop buttons if needed. Once the experiment is completed, the table’s contents become available in the results area on the right part of the experiment module. Users can read and save the original data of the experiment using the Load Data and Save Data buttons.
In the results area, the first function row contains buttons for operating on the experimental data. The first button, labeled Save Table, exports the table results to xlsx, csv, and tex files. The subsequent buttons labeled Converge control the convergence y-lim types and generate convergence plots based on the selected results from the table. The Pareto button draws the Pareto front and non-dominated solutions based on the selected results.
The edit box controls the formatting type of the table result. Following this, the drop-down box allows users to select performance metrics, and subsequent drop-down boxes determine the result display type, statistical test method, base algorithm, and result highlighting type, respectively. After selecting the statistical test method, calculated statistical results denoted by “+/-/=” followed by each metric result can also be displayed in the table. MToP offers two non-parametric methods, the Wilcoxon signed rank test and the Wilcoxon rank-sum test, for statistical tests. The base algorithm drop-down button controls the base algorithm in the statistical test, while the highlight button highlights the best and worst algorithm’s metric results for each problem or task.
@@ III-C Executing via Alternative Command Line
In addition to GUI execution, MToP offers alternative options for conducting experiments via the command line. By inputting the algorithm list, problem list, and other parameters into the script file mto.m, experiments can be executed directly from the command line. Upon completion, the experiment data MTOData.mat is returned. Similarly, by inputting the data into the metrics function, users can obtain the calculated metric results.
@@ IV Extending MToP
Users can seamlessly integrate new algorithms, problems, and performance metrics into MToP. By adhering to the code pattern outlined in Section II-C, users can develop and implement their code files with ease.
@@ IV-A Adding New Algorithms
To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the Algorithms/ folder. This class file should inherit from the base class Algorithm.m. Subsequently, users can implement the algorithmic framework following the provided code pattern, as demonstrated in Listing LABEL:list:algo. Utilizing functions such as Initialization(), Generation(), and Selection() provided by MToP is recommended. However, users can also choose to reimplement these functions to suit their specific requirements. Additionally, users seeking to create customizable parameters that can be modified via the MToP GUI should implement the getParameter() and setParameter() APIs within the new algorithm class file.